OpenStudy (anonymous):

Please could someone let me know if my answer is correct? The problem is the integral of1/(3x-2)dx with the upper limit 1 and the lower limit 0. I have the answer of: -ln(2)/(3). s this answer correct? Thank you!

OpenStudy (anonymous):

question: is the equation like this? \[\int\limits_{}^{}\frac{ dx }{ 3x - 2 }\] if so then the answer should be: \[\frac{ 1 }{ 3 } \ln 3x - 2\] because the derivative of 3x is 3dx. so you need to put 1/3 outside. also: dx/ x+1 = ln (x+1) only if the numerator is the derivative of the denominator.

OpenStudy (anonymous):

Are you a FLVS student or CCA?

OpenStudy (anonymous):

ok, so I get the part about 1/3 on outside. Because also worked it out to be the answer of: 1/3 [-ln2]. My limits changed to upper 1and lower -2 =1/3[ ln |u|] =1/3[ln1-ln2] =1/3[-ln2] w

OpenStudy (anonymous):